OpenStudy (anonymous):

find the limit as x approaches 0 for sinx/1-cosx

OpenStudy (aravindg):

well treat x=\(2 \theta\)

OpenStudy (aravindg):

then apply \(\sin 2 \theta \) and \(1-\cos 2 \theta \) formula

OpenStudy (aravindg):

cancel terms apply limits be happy !

OpenStudy (anonymous):

thank you so much!

OpenStudy (aravindg):

yw :)